Consider the following sequence of successive numbers of the 2^k-th power:
1, 2^2^k, 3^2^k, 4^2^k, 5^2^k, ...
Show that the difference between the numbers in this sequence is odd for all k ∈ N
1
Expert's answer
2020-11-18T19:10:35-0500


Let "a_m=m^{2k}, m\in \Z^+" and "a_n=n^{2k}, n\in \Z^+" be two numbers in this sequence.




"n^{2k}-m^{2k}=(n^k-m^k)(n^k+m^k)"

If "m" and "n" are not consistent numbers the difference "n^{2k}-m^{2k}" may be either odd or even.

The difference "4^{2k}-2^{2k}" is even.

The difference "4^{2k}-3^{2k}" is odd.


If "m" and "n" are not consistent numbers

Suppose "n=m+1, m\in \Z^+"

If "m" is even, then "n=m+1" is odd:



"\begin{cases}
 m &\text{is even } \\
 n &\text{is odd } 
\end{cases}=>\begin{cases}
 m^k &\text{is even } \\
 n^k &\text{is odd } 
\end{cases}=>""=>\begin{cases}
 n^k-m^k &\text{is odd } \\
 n^k+m^k &\text{is odd } 
\end{cases}=>(n^k-m^k)(n^k+m^k) \text{is odd}"

If "m" is odd, then "n=m+1" is even:



"\begin{cases}
 m &\text{is odd } \\
 n &\text{is even } 
\end{cases}=>\begin{cases}
 m^k &\text{is odd } \\
 n^k &\text{is even } 
\end{cases}=>""=>\begin{cases}
 n^k-m^k &\text{is odd } \\
 n^k+m^k &\text{is odd } 
\end{cases}=>(n^k-m^k)(n^k+m^k) \text{is odd}"

Therefore the difference between the consistent numbers in this sequence is odd for all k ∈ N.